OpenStudy (anonymous):

The difference between the digits in a two-digit numeral is 3. The units' digit is twice the tens' digit. Find the numeral.

OpenStudy (anonymous):

We have two numbers in [0,9] a,b where b=2a and b-a=3. Substitute b=2a into b-a=3 leaving 2a-a=3, so a=3; substitute a=3 into b=2a to find b=6, so our numeral is 36.

OpenStudy (anonymous):

I would suggest that it may be easiest to list some two digit numbers where the units digit is twice the tens eg 12, 24, 36, 48 and then check against the second criteria, ie a difference of three.....and the winner is..

OpenStudy (anonymous):

THANKS!